24 Ω 412 5Ω 300 Ω 100 2 200 Ω 1212 3Ω 12V 4Ω 5V 3Ω 2Ω A )3.0A Determine the current through each resistor in the circuit shown above. Determine the voltage across each resistor. Find the power dissipated in each resistorplease answer for the 2nd,3rd and 4th diagrams please.please and please only answer if you know you will answer for the 2nd 3rd and 4th diagrams please!!! this is my last question of the month please!!! I mean answer all the diagrams excluding the one on the top left
Answer #1

consider circuit 2

let current, I leave 19 V battery and i leave 12 V battery
so from kirchoff's current law, current through 200 ohm resistor = i + I

from kirchoff's voltage law
19 = 300 I + 200i + 200I = 500I + 200i
12 = 100i + 200i + 200I = 300i + 200I
i = (12 - 200I)/300

19 = 500I + (24 - 400I)/3
57 = 1500I + 24 - 400I = 1100I + 24
I = 0.03 A
i = 0.02 A

hence
current through 300 ohm resistor = 0.03 A
voltage drop = 0.03*300 = 9 V
power dissipated = 9*0.03 = 0.27 W

current through 100 ohm resistor = 0.02 A
voltage drop = 0.02*100 = 2 V
power dissipated = 2*0.02 = 0.04 W

current through 200 ohm resistor = 0.05 A
voltage drop = 0.05*200 = 10 V
power dissipated = 10*0.05 = 0.5 W